ASYMPTOTIC GEOMETRY AND CONFORMAL TYPES OF CARNOT–CARATHÉODORY SPACES

VLADIMIR A. ZORICH

Abstract. An intrinsic definition in terms of conformal capacity is proposed for the conformal type of a Carnot–Carathéodory space (parabolic or hyperbolic). Geometric criteria of conformal type are presented. They are closely related to the asymptotic geometry of the space at infinity and expressed in terms of the isoperimetric function and the growth of the area of geodesic spheres. In particular, it is proved that a sub-Riemannian manifold admits a conformal change of metric that makes it into a complete manifold of finite volume if and only if the manifold is of conformally parabolic type. Further applications are discussed, such as the relation between local and global invertibility properties of quasiconformal immersions (the global homeomorphism theorem).